Dickinsen Company gathered the following data for December:

Planned Actual
Sales price per unit $5.80 $6.00
Number of units of sales × 820,000 × 805,000
Total sales $4,756,000 $4,830,000

Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Compute the revenue price variance.

$

b. Compute the revenue volume variance.

$

c. Compute the total revenue variance.

$
Answer #1

a) Revenue price variance = (5.80-6)*805000 = -161000 F

b) Revenue volume variance = (820000-805000)*5.8 = 87000 U

c) Total revenue variance = -74000 F